# Learning to Generate Unit Tests for Automated Debugging

### v1

@@ Abstract
Unit tests (UTs) play an instrumental role in assessing code correctness as well as providing feedback to a large language model (LLM) as it iteratively debugs faulty code, motivating automated test generation. However, we uncover a trade-off between generating unit test inputs that reveal errors when given a faulty code and correctly predicting the unit test output without access to the gold solution. To address this trade-off, we propose UTGen, which teaches LLMs to generate unit test inputs that reveal errors along with their correct expected outputs based on task descriptions and candidate code. We integrate UTGen into UTDebug, a robust debugging pipeline that uses generated tests to help LLMs debug effectively. Since model-generated tests can provide noisy signals (e.g., from incorrectly predicted outputs), UTDebug (i) scales UTGen via test-time compute to improve UT output prediction, and (ii) validates and back-tracks edits based on multiple generated UTs to avoid overfitting. We show that UTGen outperforms UT generation baselines by 7.59% based on a metric measuring the presence of both error-revealing UT inputs and correct UT outputs. When used with UTDebug, we find that feedback from UTGen’s unit tests improves pass@1 accuracy of Qwen-2.5 7B on HumanEvalFix and our own harder debugging split of MBPP+ by over 3% and 12.35% (respectively) over other LLM-based UT generation baselines.
@@ 1 Introduction
With rapid advancements in training large language models (LLMs; Achiam et al., 2023; Anthropic, 2024; Gemini et al., 2023; Dubey et al., 2024), enhancing their coding abilities has garnered significant attention (Chen et al., 2021; Li et al., 2023; Roziere et al., 2023; Guo et al., 2024b; Hui et al., 2024, inter alia). However, these models are far from perfect and – much like human-written code – model-written code contains errors. Human developers often improve their code through test-driven development with iterative debugging, i.e., identifying failure cases by providing example inputs and their expected outputs, reasoning over causes of failure, and modifying the code to address the issues. Unit tests (UTs) are one form of such pairings: code inputs and expected outputs that test individual functionalities in a piece of code. A failing unit test provides targeted feedback, not only identifying that a piece of code is not operating as expected but also helping to localize the error to one part of the code (Maximilien & Williams, 2003; Nagappan et al., 2008; Ficco et al., 2011). Models have similarly been shown to benefit from iterative debugging based on explicit or implicit feedback stemming from failing unit tests (Chen et al., 2023b; Moon et al., 2023). However, the ability to provide feedback and debug incorrect code is often bottlenecked by the availability of (failing) unit tests for a given problem.
While several coding benchmarks come with human-written UTs for evaluation purposes, these suites of UTs are typically small due to the laborious annotation process (Liu et al., 2024). Unit test collection is challenging for two reasons: first, it requires sampling inputs that are likely to trigger an error. For example in Figure 1, the unit test: would not trigger any error (despite the fact that the code is incorrect since the error happens to not be triggered for multiples of 10), while another unit test: would lead to an incorrect output, revealing a bug in the function. Secondly, UTs require expected outputs, i.e., the desired behavior of the code being tested must be known. For example, in Figure 1, the expected output 131 must be given in advance. Due to these challenges, prior work employs LLMs to generate unit test inputs at random (i.e., without conditioning on faulty code to generate failing inputs) and often uses the gold (correct) code solution for generating outputs (Chen et al., 2023a; Chae et al., 2024), which is generally not available at test time. Therefore, these approaches do not scale to automated debugging, which requires an online and automated method for dynamically generating UTs as problems are iteratively debugged by the model. To address this gap, we pose the following research questions:
@@
RQ2: How well do LLMs perform zero-shot UT generation, and how can we improve their abilities?
@@
RQ3: How do we best use an automated but potentially noisy UT generator to improve code debugging?
@@
To address RQ1, we characterize two desirable properties of unit test generators (in Section 3.1): 1) high attack rate, i.e., given a faulty code, the unit test generator should generate UT inputs that are likely to trigger errors;111Here, an ‘error’ means either a syntax error, a runtime error, or an assertion error where the result of the target code does not match the expected output of the UT. We expand on this in Section 3. 2) high output accuracy, ensuring that the corresponding output of the unit test is consistent with the task description (and that of a correct solution). For instance, in Figure 1, would lead to a lower attack rate, as it does not trigger any errors, while the generated UT does trigger the error. While one test is attacking and another is not, both have high output accuracy, since in both cases the output is correct. By design, we can measure these properties via three intrinsic metrics: measuring attack rate and output accuracy independently and then, crucially, how often a UT generator can generate both failing inputs along with correct UT outputs based only on the problem description and a faulty code, i.e., without access to the gold solution. We benchmark the ability of various LLMs to act as zero-shot unit test generators using attack rate and output accuracy (i.e., intrinsic metrics for UT generation) and show that coding LLMs generally struggle with this task (cf. Section 5.1).
@@
Moreover, addressing RQ2, we find that, without any training, zero-shot models often exhibit a strong trade-off between attack rate and output accuracy, since the UTs that are most likely to trigger errors (i.e., higher attack-rate UTs) are generally more challenging edge cases, making it harder for the model to predict their output (i.e., lower output accuracy). To address this trade-off, we propose training models; due to a lack of dedicated datasets for unit test generation in prior work, we introduce UTGen, a data creation and training recipe (Section 3.2) that bootstraps training data for UT generation from existing code generation benchmarks by perturbing code to simulate errors, generating failing unit test and augmenting it with chain-of-thought-style rationales (Wei et al., 2022). We show that training via UTGen better balances this trade-off and yields a higher number of unit tests that have both an input that triggers an error and also predicts the correct output by 7.59% more over merely prompting models to generate failing UTs (in Section 5.1).
@@
Building on these intrinsic evaluations, we examine RQ3, testing our generated UTs on a downstream, i.e., extrinsic task: code debugging. Here, automated UT generation methods face additional challenges: unlike human-generated gold UTs – which have correct outputs but require human involvement – generated UTs provide noisy feedback, as the UT might fail to reveal the buggy code’s errors or have an incorrectly predicted output. To mitigate this issue, we propose UTDebug, an improved multi-turn debugging method that addresses these challenges in two major ways (cf. Section 3.3). First, we improve the output accuracy of generated UTs by scaling test-time compute via self-consistency (Wang et al., 2022), taking a vote over multiple output samples to obtain better results. Moreover, to prevent overfitting to noisy feedback from inaccurate UTs, we regularize the debugging process by generating multiple UTs and accepting code edits based on their effect on a validation suite of UTs. Specifically, if a proposed change results in fewer generated tests passing (i.e., hurts the code quality) we allow the debugger to backtrack to a previous stage and re-attempt debugging. We then plug UTs generated from multiple LLM-based methods – including UTGen – into UTDebug to extrinsically to measure the utility of generated UTs as feedback (Chen et al., 2023b; Zhong et al., 2024), as illustrated by the final output of debugging in Figure 1. Our results show that on our most challenging subset of MBPP+ problems, UTDebug with UTs generated by UTGen improves pass@1 accuracy of Qwen2.5 by (absolute) compared to debugging without unit tests, and by over a zero-shot baseline.
Our key contributions are:
We bring attention to the trade-off between attack rate and output accuracy, releasing a challenging dataset with subtle errors that require models to balance these properties. We find that this data is challenging even for models trained extensively on code.
We introduce UTGen for training UT generators that alleviates the trade-off between attack rate and output accuracy through training, building UT generators with improved performance on both.
Building on the scalability of UTGen, we introduce UTDebug, a debugging method that uses test-time scaling and a back-tracking validation strategy enabled by a suite of generated unit tests to improve debugging performance and avoid overfitting.
@@ 2 Related Work
Automatic Unit Test Generation.
While unit tests play a pivotal role in assessing code correctness, manually writing a suite of unit tests is laborious and often infeasible (Chen et al., 2023a; Liu et al., 2024). Therefore, prior work in software testing has often focused on automatic unit test generation (King, 1976; Cadar et al., 2008; Holler et al., 2012; Cha et al., 2015, inter alia). With the application of LLMs to program synthesis, recent work has explored using LLMs for automatic unit test generation (Chen et al., 2023a; Schäfer et al., 2023; Liu et al., 2024). Notably, Schäfer et al. (2023) and Liu et al. (2024) focus solely on unit test input generation, relying on the gold (correctly implemented) function being tested to generate the corresponding output of the unit test. Both prompt strong frontier models such as GPT-4, followed by additional input mutations (Liu et al., 2024) or iteratively prompting the LLM (Schäfer et al., 2023). In contrast, our open-source models trained using UTGen automatically generate both input-output pairs of unit tests, i.e., at inference time, UTGen generates a candidate input and the corresponding output as per the task description without access to the gold solution. While Chen et al. (2023a) also generate both input-output unit test pairs, they only rely on standard LLM prompting to produce unit tests as a means of improving code-generation performance. As a result, their method does not evaluate or optimize for the validity of generated UTs, especially output accuracy; on the other hand, in UTGen, we explicitly model output accuracy and train the models to reason about the output of the unit test correctly.
LLM Debugging. Using LLMs for debugging faulty code (also known as program repair) has been a subject of prior work. Debugging approaches can be categorized into two groups: those training models to debug (Moon et al., 2023; Ni et al., 2024; Chae et al., 2024) and those providing external feedback to pretrained models (Chen et al., 2023b; Olausson et al., 2023; Zhong et al., 2024). Both groups share a reliance on (gold) unit tests to train models and/or provide feedback. Thus, UTGen is complementary to both kinds of debugging methods and can be easily integrated into either by using unit tests generated by UTGen when human-written unit tests are scarce or unavailable. To this end, in Section 5 we show that UTGen’s generated unit tests can be used to provide feedback to LLMs via rubber-duck debugging proposed in Chen et al. (2023b). Additionally, in Section 3.3, we propose UTDebug, an improvised debugging pipeline that handles noisy feedback from inaccurately generated UTs via test-time scaling and backtracking. Orthogonal to our work, another direction of prior work involves curating benchmarks to measure LLMs’ debugging abilities (Muennighoff et al., 2024; Jain et al., 2024; Tian et al., 2024) – which we also use and build upon in our experiments to showcase the effectiveness of UTGen.
@@ 3.1 Background and Task Setup
Given a natural language task description for a coding task, we focus on generating unit tests (UTs) in the format of input-output pairs that are consistent with the task description . Our setup is consistent with Chen et al. (2023a) and Jain et al. (2024) who also consider unit tests in the form of input-output pairs generated without utilizing the correct implementation of the function. More generally, our setting is akin to parameterized unit testing (Tillmann et al., 2010) and uses the notion of functional correctness, i.e., measuring correctness by simulating an exhaustive set of scenarios (UT inputs) and ensuring that the function performs as expected (correct output as per the problem description ).222We note that the correctness of a code or function can also be determined via formal verification, i.e., by mathematically proving it functions correctly in all situations. However, we do not consider this paradigm in our work, as we focus on empirical methods and applications, e.g., code debugging.
Notation and Desiderata. Let denote the natural language description of the function to be implemented (top yellow box in Figure 1). We assume that this description specifies the input space and output space . Furthermore, let the set of all functions that correctly solve the task be . Then, a valid unit test for task is:
@@
For example, in Figure 1, 120 is a valid input, as it is a number, whereas "apple" is not. Similarly, 121 is the expected output of the function (given 120 as input), while 122 would be an invalid output. Thus, (120, 121) is a valid unit test for the task.
Unit Test Generator. Addressing RQ1, we define the desirable properties of an automatic unit test generator , parameterized by . Ideally, should generate valid unit tests from a task description , i.e., without any manual intervention. However, to account for downstream utility of unit tests, we denote a potentially buggy code under testing as . If is faulty or implemented incorrectly, it should fail one or more valid unit tests generated by . Therefore, given a buggy piece of code for the problem , a desirable unit test generator should be able to efficiently generate failing unit tests: , such that , i.e., a valid unit test input that uncovers the errors in . Moreover, must also predict the correct UT output , i.e., the output of the correct code. We test both generators of the form that only consider the description, as well as debugging-style generators of the form that start with a buggy code solution and take it into account in generating unit tests for it. Empirically, we find that generators of the form lack sufficient context to generate tests that help reveal errors, limiting their utility for debugging. We therefore focus on training functions of the form .
3.2 UTGen: Training LLMs for Unit Test Generation
While several lines of prior work focus on curating training data for improving code generation (Guo et al., 2024a; Muennighoff et al., 2023), there is a general lack of dedicated datasets for training the desired UT generator outlined above. Therefore, to improve along the lines of RQ2, we design a training recipe that bootstraps this data from training datasets for code generation i.e., a collection of problem descriptions and their corresponding code implementations. After collecting these examples, we perturb their gold solutions to generate faulty code for which we will be generating unit tests. For each example, we sample multiple unit tests and select the ones that are adversarial w.r.t. to the buggy code (i.e., inputs that cause it to fail), which we use to construct our training data. Below we describe each step:
@@ 3.2 UTGen: Training LLMs for Unit Test Generation
We start with a collection of coding problems with problem descriptions and gold codes. Specifically, we use publicly available data from Tulu-3 (Lambert et al., 2024) due to its large scale and the improvements noted by Lambert et al. (2024) when finetuning off-the-shelf LLMs on coding tasks. We filter it to focus on Python code with functional abstractions (further details in Appendix B). This yields a total of 48.3K unique code problems. Corresponding to each problem description (prompt), the dataset contains a correct target code solution intended for finetuning, which we use as the reference (gold) code solution () as illustrated in Figure 2 (I).333The code solutions in the Tulu-3 SFT mix have either been written by humans or by frontier models and are thus highly likely to be correct (Lambert et al., 2024). However, in order to train an adversarial unit test generator we would need access to incorrect or faulty code solutions that can be debugged, as unit tests must have some error to attack. We obtain these by using the LLM to perturb the reference code solution in realistic ways (prompt and details given in appendix Appendix E).
Annotating Unit Tests. As mentioned in Section 3.1, one of the goals of the unit test generator is to be able to not only generate valid unit tests, but failing unit tests (that trigger errors). To facilitate this, given a problem description, reference code , and buggy candidate code , we sample different unit test inputs (via the prompts in Appendix E). A unit test is failing if , i.e., if the output of the candidate fails to match the reference output (cf. Figure 2 (II)). Note that, while the LLM can be used to generate the output of unit test, it can often be inaccurate (Jain et al., 2024; Gu et al., 2024); so to ensure output accuracy, we use the output of the reference code during training. We revisit the discussion of how accurate LLMs are at output prediction during inference in Section 5.
@@
When training LLMs according to UTGen, the input to the LLM is the same prompt used for sampling unit tests (listed in Appendix E) and the output is an adversarial unit test. The goal here is to improve both output accuracy and attack rate jointly via supervised finetuning using the negative log-likelihood loss. Since LLMs have been known to struggle at the task of output prediction (Jain et al., 2024), we train the model to generate chain-of-thought (CoT; Wei et al., 2022) reasoning before predicting the UT output. To this end, we employ the post-hoc rationalization procedure outlined in Zelikman et al. (2022) – given the entire UT , we ask the LLM to generate rationales supporting why is the output corresponding to the input . Then, to create the supervision data, we add these rationales as CoTs prior to the output prediction (see Figure 2 (III)).
3.3 UTDebug: Debugging with Generated Unit Tests

### v2

@@ Abstract
Unit tests (UTs) play an instrumental role in assessing code correctness as well as providing feedback to large language models (LLMs), motivating automated test generation. However, we uncover a trade-off between generating unit test inputs that reveal errors when given a faulty code and correctly predicting the unit test output without access to the gold solution. To address this trade-off, we propose UTGen, which teaches LLMs to generate unit test inputs that reveal errors along with their correct expected outputs based on task descriptions. Since model-generated tests can provide noisy signals (e.g., from incorrectly predicted outputs), we propose UTDebug that (i) scales UTGen via test-time compute to improve UT output prediction, and (ii) validates and backtracks edits based on multiple generated UTs to avoid overfitting, and helps LLMs debug effectively. We show that UTGen outperforms other LLM-based baselines by 7.59% based on a metric measuring the presence of both error-revealing UT inputs and correct UT outputs. When used with UTDebug, we find that feedback from UTGen’s unit tests improves pass@1 accuracy of Qwen2.5 32B on HumanEvalFix and our own harder debugging split of MBPP+ by over 3.17% and 12.35% (respectively) over other LLM-based UT generation baselines. Lastly, we demonstrate that UTGen is a better judge for code correctness, outperforming a state-of-the-art trained 8B reward model by 4.43% on HumanEval+ with best-of-10 sampling using Qwen2.5 7B.111Code: https://github.com/archiki/UTGenDebug
@@
Learning to Generate Unit Tests for Automated Debugging
Archiki Prasad Elias Stengel-Eskin Justin Chih-Yao Chen Zaid Khan Mohit Bansal UNC Chapel Hill
@@ 1 Introduction
With rapid advancements in training large language models (LLMs; Achiam et al., 2023; Anthropic, 2024; Gemini et al., 2023), enhancing their coding abilities has garnered significant attention (Chen et al., 2021; Li et al., 2023; Roziere et al., 2023; Guo et al., 2024b, inter alia). However, these models are far from perfect and – much like human-written code – model-written code contains errors. Human developers often improve their code through test-driven development, i.e., identifying failure cases by providing example inputs and their expected outputs – referred to as unit tests (UTs) – to test individual functionalities, reasoning over causes of failure, and modifying the code to address the issues (Maximilien and Williams, 2003; Nagappan et al., 2008; Ficco et al., 2011; Beck, 2022). Models have similarly been shown to benefit from iterative debugging based on explicit or implicit feedback stemming from failing unit tests Chen et al. (2023b); Moon et al. (2023). However, the ability to provide feedback and debug incorrect code is often bottlenecked by the availability of (failing) unit tests for a given problem.
While several coding benchmarks come with human-written UTs for evaluation purposes, these suites of UTs are typically small due to the laborious annotation process Liu et al. (2024b). Unit test collection is challenging for two reasons: first, it requires sampling inputs that are likely to trigger an error. For example in Fig. 1, the unit test: (120)==121 would not trigger any error (despite the fact that the code is incorrect for non-multiples of 10), while another unit test: (123)==131 would lead to an incorrect output, revealing a bug in the function. Secondly, UTs require expected outputs (e.g., 131 in the previous UT), i.e., the desired behavior of the code being tested must be known. Due to these challenges, prior work employs LLMs to generate unit test inputs at random (i.e., without conditioning on faulty code to generate failing inputs) and often uses the gold (correct) code solution for generating outputs Chen et al. (2023a); Chae et al. (2024). Therefore, these approaches do not scale well at inference time, which requires an online and automated method for generating UTs as the LLMs solve problems on the fly. To bridge this gap, we pose the following research questions:
@@
RQ2: How well do LLMs perform zero-shot UT generation, and how can we improve them?
@@
RQ3: How do we best use an automated but potentially noisy UT generator to improve code debugging and generation?
@@
To address RQ1, we characterize two desirable properties of unit test generators (in Sec. 3.1): 1) high attack rate, i.e., given a faulty code, the unit test generator should generate inputs that are likely to trigger errors;222Here, an ‘error’ means either a syntax/runtime error, or when the result of the target code does not match the expected output of the UT. We expand on this in Sec. 3. 2) high output accuracy, ensuring that the UT output is consistent with the task description (and that of a correct solution). For instance, in Fig. 1, (120) would lead to a lower attack rate, as it does not trigger any errors, while (123)==131 does (high attack rate); however, both have high output accuracy as in both cases the output is correct. We measure these properties via three intrinsic metrics: measuring attack rate and output accuracy independently and then, crucially, how often a UT generator can generate both failing inputs along with correct UT outputs based only on the problem description and a faulty code. We benchmark the ability of LLMs to act as zero-shot unit test generators and show that coding LLMs generally struggle (cf. Sec. 5.1).
@@
Moreover, addressing RQ2, we find that, without finetuning, zero-shot models often exhibit a strong trade-off between attack rate and output accuracy. In other words, the UTs that are most likely to trigger errors (i.e., higher attack rate) are generally more challenging edge cases, making it harder for the model to predict their output (i.e., lower output accuracy). Due to a lack of dedicated datasets for unit test generation in prior work, we introduce UTGen, a data creation and training recipe (Sec. 3.2) that bootstraps training data for UT generation from existing code generation datasets by perturbing code to simulate errors, generating failing unit test and augmenting it with chain-of-thought rationales Wei et al. (2022). We show that training via UTGen better balances this trade-off and yields a higher number of unit tests that have both attacking inputs and correct outputs, with 7.59% absolute improvement over prompting models to generate failing UTs.
@@
Finally, we examine RQ3, utilizing our generated UTs on downstream, i.e., extrinsic tasks: code debugging and code generation. Here, automated UT generation methods face additional challenges: unlike human-generated gold UTs – which have correct outputs but require human involvement – generated UTs provide noisy feedback, as the UT might fail to reveal the buggy code’s errors or have an incorrectly predicted output. To mitigate this issue, we propose UTDebug, an improved multi-turn debugging method that addresses these challenges in two ways (cf. Sec. 3.3): (i) we improve the output accuracy of generated UTs by scaling test-time compute via self-consistency (Wang et al., 2022); (ii) we regularize the debugging process by generating multiple UTs and accepting code edits only if the revised code passes more generated UTs, backtracking edits otherwise. We then plug UTs generated from multiple LLM-based methods into UTDebug in Sec. 5.2, finding that on our most challenging subset of MBPP+ problems, UTDebug with UTs generated by UTGen improves pass@1 accuracy of Qwen2.5 7B by 12.9% (absolute) compared to debugging without UTs, and by 5.3% over a zero-shot unit test generation baseline. We demonstrate that by generating UTs with UTGen and computing the rate of passing UTs, we can better judge code correctness than using scores from trained, state-of-the-art 8B reward model (RM). For best-of-10 sampling with Qwen2.5 7B, UTGen improves accuracy by 4.43% over the RM on HumanEval+.
@@ 2 Related Work
Automatic Unit Test Generation. Manually writing unit tests is laborious and often infeasible Chen et al. (2023a); Liu et al. (2024b). Consequently, past research explores automatic UT generation (King, 1976; Cadar et al., 2008; Holler et al., 2012; Cha et al., 2015, inter alia). The advent of LLMs has spurred recent efforts in using them for UT generation Chen et al. (2023a); Schäfer et al. (2023); Liu et al. (2024b). Specifically, Schäfer et al. (2023) and Liu et al. (2024b) focus on generating unit test inputs via prompting frontier LLMs like GPT-4 and/or iterative prompting, assuming access to the gold solution. In contrast, our models, trained with UTGen, generate both input-output UT pairs based on the task description without relying on the gold implementation. While Chen et al. (2023a) also generate input-output UT pairs using standard LLM prompting, their primary focus is code generation – not the quality of generated UTs. On the other hand, we directly model the desiderata of UTs including output accuracy, and demonstrate its utility on code generation and debugging.
LLM Debugging. Using LLMs for debugging faulty code, or program repair, has been extensively studied. Debugging approaches are divided into those training models to debug Moon et al. (2023); Ni et al. (2024); Chae et al. (2024) and those providing external feedback to pretrained models Chen et al. (2023b); Olausson et al. (2023); Zhong et al. (2024). Both rely on (gold) unit tests for training or feedback. Thus, UTGen complements both methods by providing generated unit tests when human-written tests are scarce or unavailable. In Sec. 3.3, we introduce UTDebug, a debugging pipeline that addresses noisy feedback from inaccurate unit tests through test-time scaling and backtracking. Moreover, in Sec. 5 we show that UTGen’s unit tests can effectively provide feedback to LLMs for code generation and debugging.
@@ 3.1 Background and Task Setup
Given a natural language task description for a coding task, we focus on generating unit tests (UTs) in the format of input-output pairs that are consistent with the task description . Our setup is consistent with Chen et al. (2023a) and Jain et al. (2024) who also consider unit tests in the form of input-output pairs generated without utilizing the correct implementation of the function. More generally, our setting is akin to parameterized unit testing Tillmann et al. (2010) and uses the notion of functional correctness, i.e., measuring correctness by simulating an exhaustive set of scenarios (UT inputs) and ensuring that the function performs as expected (as per the problem description ).
Notation and Desiderata.
Let denote the natural language description of the function to be implemented (top yellow box in Fig. 1). We assume that this description specifies the input space and output space . Furthermore, let the set of all functions that correctly solve the task be . Then, a valid unit test for task is:
@@
For example, in Fig. 1, 120 is a valid input, as it is a number, whereas "apple" is not. Similarly, 121 is the expected output of the function (given 120 as input), while 122 would be an invalid output. Thus, (120, 121) is a valid unit test for the task.
Unit Test Generator.
Addressing RQ1, we define the desirable properties of an automatic unit test generator , parameterized by . Ideally, should generate valid unit tests from a task description , i.e., without any manual intervention. However, to account for downstream utility of unit tests, we denote a potentially buggy code under testing as . If is faulty or implemented incorrectly, a desirable unit test generator should be able to efficiently generate failing unit tests: , such that , i.e., a valid unit test input that uncovers the errors in . Moreover, must also predict the correct UT output , i.e., the output of the correct code. We study generators of the form , which consider only the description, and debugging-style generators of the form , which also consider a buggy code solution . Empirically, we find that lacks sufficient context to generate effective error-revealing tests. Therefore, we focus on training UT generators of the form .
3.2 Training LLMs for Unit Test Generation
While prior work focus on curating training data for improving code generation Guo et al. (2024a); Muennighoff et al. (2023), there is a general lack of dedicated datasets for training the desired UT generator outlined above. Therefore, to improve along the lines of RQ2, we design, UTGen, a training recipe that bootstraps this data from training datasets for code generation i.e., a collection of problem descriptions and their corresponding code implementations.
@@ 3.2 Training LLMs for Unit Test Generation
We start with a collection of coding problems with problem descriptions () and gold codes () as illustrated in Fig. 2 (I). Specifically, we use publicly available data from Tulu-3 Lambert et al. (2024a) due to its large scale and the improvements noted by Lambert et al. (2024a) when finetuning LLMs on coding tasks.333The code solutions in the Tulu-3 SFT mix have either been written by humans or by frontier models and are thus highly likely to be correct Lambert et al. (2024a). We filter it to focus on Python code with functional abstractions (further details in Appendix B). However, in order to train an adversarial unit test generator we would need access to incorrect or faulty code solutions that can be debugged, as unit tests must have some error to attack. We obtain these by using the LLM to perturb the reference code solution.
Annotating Unit Tests.
As mentioned in Sec. 3.1, one of the goals of the unit test generator is to be able to not only generate valid unit tests, but failing unit tests (that trigger errors). To facilitate this, given a problem description, reference code , and buggy candidate code , we sample different unit test inputs (via the prompts in Appendix E). A unit test is failing if , i.e., if the output of the candidate fails to match the reference output (cf. Fig. 2 (II)). Note that, while the LLM can be used to generate the output of unit test, it can often be inaccurate Jain et al. (2024); Gu et al. (2024); so to ensure output accuracy, we use the output of the reference code during training. We revisit the discussion of how accurate LLMs are at output prediction during inference in Sec. 5.
@@
When training LLMs, the input to the LLM is the same prompt used for sampling unit tests (listed in Appendix E) and the output is an adversarial unit test. The goal here is to improve both output accuracy and attack rate jointly via supervised finetuning using the negative log-likelihood loss. Furthermore, we train the model to generate chain-of-thought (CoT; Wei et al., 2022) reasoning before predicting the UT output. To this end, we employ the post-hoc rationalization procedure outlined in Zelikman et al. (2022) – given the entire UT , we ask the LLM to generate rationales supporting why is the output corresponding to the input . Then, to create the supervision data, we add these rationales as CoTs prior to the output prediction (see Fig. 2 (III)).
3.3 Generated Unit Tests for Debugging